# Pseudopotentials for Orbital-Free DFT: Capturing Nonlocality and Correcting Functional Approximants

### v1

@@ I Introduction and Background
In OF-DFT, the noninteracting kinetic energy, $T_{s}$, is evaluated approximately by a pure density functional, which we call henceforth the kinetic energy density functional or KEDF. The accuracy of OF-DFT depends on the KEDF $T_{s}[\rho]$, with magnitude equal to the total energy, for highly inhomogeneous electron densities the devolopment of KEDFs is still a challenging task, however there are good functionals sunch as LMGP[mi2019orbital, mi2019nonlocal], Huang-Carter (HC)[huang2010nonlocal], and revHC[shao2021revised] which provides one of the most accurate descriptions of metallic systems currently available. Using KEDFs allows the energy minimization of the functional in Eq. (1) to be based solely on the electron density, completely bypassing the KS orbitals, and generally leading to algorithms of linear scaling complexity. [mi2023orbital].
@@
The guiding principles to build LPPs have been to employ them in a KS-DFT calculation (i.e., with exact $T_{s}$) imposing it to reproduce the electron density, lattice constant and even the band structure of solids [mi2016first, huang2008transferable, zhou2004transferable]. For those elements for which this is possible, it is said LPPs can be employed in OF-DFT calculations. It is certainly surprising that band structures can be reproduced in pseudopotential calculations having only a local part. However, in many cases it was shown to be possible [starkloff1977local]. It turns out that LPPs can be successfully constructed for the main-group elements mentioned above and also for a large array of transition metals (provided that the semicore states are excluded from the pseudization [chi2024high]). The most accurate LPPs available to date for transition metal elements are the recently developed highly accurate LPPs by the Huang group [chi2024high]. The valence density of transition metals consists of both localized $d$ electrons and delocalized sspp conduction electrons, a regime in which most existing KEDFs are constructed from the response of homogeneous or nearly homogeneous electron gas struggle to provide accurate kinetic energies. In addition, reliable LPPs must account for semicore $s$-like states that lie close in energy to the valence $d$ bands and influence bonding and screening. Since OF-DFT approximates $T_{s}[\rho]$ strictly as a functional of the electron density, it lacks explicit angular-momentum dependence and therefore cannot distinguish $d$-state character from $s$ or $p$ contributions. These combined factors explain why transition metals are particularly demanding within OF-DFT and motivate the construction of pseudopotentials specifically designed to mitigate these limitations. In this work, semicore states are not explicitly included in the optimization procedure to preserve numerical stability across the periodic table; however the proposed methodology corrects the KEDF and the PP at the same time by reproducing correct scattering properties and remain transferable in bulk and cluster environments.
@@ II.1 Nonlocal Pseudopotential Representations
Hohenberg and Kohn theorems for nonlocal external potentials due to Gilbert [gilbert1975hohenberg] apply and guarantee the existence of the above map, even though it is generally not bijective. The non-bijectivity is due to the fact that for a given choice of $\hat{v}_{nl}$ one can find an associated $v_{loc}({\mathbf{r}})$ such that the KS equations yield the target density $n({\mathbf{r}})$. Therefore there are multiple (potentially an infinite number) of $\hat{v}_{NLPP}$ that have nonzero nonlocal part and that yield the same electron density, $n({\mathbf{r}})$. We call the set of such NLPPs a NLPP representation. Typically, there is one NLPP having the property that beyond $R_{cut}^{nl}$ not only the pseudodensity matches the all-electron density, but also the pseudo KS wavefunctions match the all-electron KS wavefunction for the isolated atom [kleinman1982efficacious].
@@ III Simultaneously approximating $\Delta v_{LPP}({\mathbf{r}})$ and the OF-DFT OEP problem
This definition clearly shows two goals for $\Delta v_{LPP}({\mathbf{r}})$: (1) it includes the effect of the nonlocal part of the KS-DFT PP; (2) it corrects for the fact that KEDFs are approximate, making sure that the density of the OF-DFT system matches the one of the KS-DFT system. Hence, the selection of the KEDF is based on computationally efficent functionals shuch as TF+vW. Thus, the result of approximating $\Delta v_{LPP}({\mathbf{r}})$ is to simultaneously achieve the two goals.
@@
Important aspects to note are that $\Delta v_{LPP}({\mathbf{r}})$ is: (1) short ranged around each ion, as it is only nonzero for $0\leq|{\mathbf{r}}-\mathbf{R}_{I}|\leq r_{cut}$; (2) of spherical symmetry around each ion. Thus, the resulting LPP, given by $v_{loc}^{\hat{v}_{nl}}(r)+\Delta v_{LPP}(r)$ maintains the correct Coulomb decay in agreement with the target NLPP. Finally, we note that there is freedom in choosing the cutoff radii ($r_{cut}$). We use the value of the largest of the $R_{cut}^{nl}$ of the target NLPP, see supplementary Table S1. We will see in the results section that this choice might not be the most optimal given the restrictive, polynomial ansatz for $\Delta v_{LPP}(r)$. The choice of cutoff radius will deserve further consideration in future works.
Regarding the choice of polynomial ansatz for $\Delta v_{LPP}(r)$, after testing, we found that a 5th order provides a good balance between accuracy of teh LPPs and computational cost of the LPP optimizer. Reducing the order of the polynomial would result is significantly lower quality LPPs, and increasing the order would result is significant computational cost and convergence complexity for the optimizer.
@@ IV Computational Details
All calculations employed the Local Density Approximation (LDA) xc functional proposed by Perdew and Zunger [perdew1981self]. Therefore, target results in this work are given by pseudopotential KS-DFT simulations employing the LDA functional. We also tested our pseudization procedure using PBE [perdew1996generalized] for Ag recovering a similarly accurate pseudization as LDA (see Results section). Testing other xc functionals will be matter of future work.
@@ V.1 Generation of the new LPPs
where $N_{v}$ is the number of valence electrons. We note that $\Delta n\leq N_{v}$, thus $\Delta_{v}\leq 100$. In the Supplementary material (Figure S1) we show the spatial distribution of the electron density difference between the OF and KS in a representative cut plane of the most stable phases for Ag, Cu, Pd, and Mo, to validate the convergence criteria in Eq. (20).
@@ V.2 Equation of State
From the previous analysis, we have chosen to further investigate PGBRV0.2 for a selected set of atoms such as Cu, Ag, and Pd for which PGBRV0.2 is expected to improve upon the local part of the GBRV PP (i.e., the $v_{loc}^{\hat{v}_{nl}}$ mentioned before). We also present Mo for which we predict the PGBRV0.2 will not provide improved electronic structure compared to the local part of the GBRV NLPP. The analysis involves computation of EOS parameters (equilibrium volume, $V_{0}$; equilibrium energy, $E_{0}$; and bulk modulus, $B_{0}$) for FCC, BCC, and HCP phases. We compare the PGBRV0.2 results with the local part of the GBRV NLPP and HQLPP. All calculations are carried out with TF0.2vW and the nonlocal revHC KEDFs. The results are collected in Tables 1 for Ag, 2 for Cu, 3 for Pd, and 4 for Mo, where we show the relative error of $V_{0}$ and $B_{0}$ and the energy difference compared to the most stable phase of the presented elements. The EOS parameters from KS (LDA) are compared to experimental values as shown in Table S9 and show that the KS-DFT benchmark results are close to experimental ground truth for lattice constants and within 30% for bulk moduli.
@@
Even though the results presented in Tables 14, as well as in supplementary Tables S2-S5, are not quantitatively accurate, they are nonetheless reasonable. In Table S6 we show an statistical summary of Tables S2-S5, where the Mean Absolute Relative Error (MARE) values ($\approx$ 50-60%) obtained for $V_{0}$ and $B_{0}$ using the original GBRV and PSL pseudopotentials indicate high discrepancies between the OF-DFT and KS-DFT frameworks. In contrast, the new local pseudopotentials (PGBRV0.2 and PPSL0.2) significantly reduce the MARE for $V_{0}$ to 28–37%, corresponding to a significant improvement in accuracy relative to emplying the local part of the KS-DFT pseudopotentials. Although the $B_{0}$ values remain less accurate (MARE $\approx$55–60%), all materials are predicted to be bound. This results are nontrivial for OF-DFT calculations[rios2024effective, witt2021random], and as previously noted, the OF-DFT calculations employing the newly developed LPPs successfully capture several trends observed in the KS-DFT EOS parameters. To rationalize this outcome, we emphasize that the PGBRV0.2 LPP is designed not only to partially recover effects associated with the nonlocal component of the NLPP, but also, crucially, to locally invert the KS density in regions near the ionic cores. Consequently, using the PGBRV0.2 pseudopotentials yields electron densities of substantially higher quality than would be obtained using a generic, KEDF-agnostic LPP. A clear demonstration of this point is provided by the performance comparison with HQLPP pseudopotentials, which yield lower accuracy than PGBRV0.2 a more detailed quantitative comparison between PGBRV0.2 and HQLPPs performance is provided in Tables S7 of the Supplementary Information. Upon switching to the revHC functional (Table S8), the PGBRV0.2 potentials maintain comparable accuracy (MARE $\approx$ 20–40%), while the HQLPPs exhibit larger deviations. These results confirm that the new pseudopotentials achieve improved consistency with KS-DFT and remain transferable across different KEDFs. This discrepancy does not reflect a flaw in the HQLPP pseudopotentials, as they were explicitly constructed to best represent electron densities calculated with the exact KEDF.
@@
FCC
KS GBRV 0.00 -4023.32 0.00
TF0.2vW GBRV 43.0 -3719.80 44.6
PGBRV0.2 20.1 -4546.42 22.3
HQLPP 34.7 -4054.91 50.4
revHC GBRV 58.1 -3731.58 56.8
PGBRV0.2 13.3 -4380.85 12.2
HQLPP 22.6 -3998.92 50.4
HCP
KS GBRV 0.00 0.01 0.00
TF0.2vW GBRV 35.8 0.14 59.5
PGBRV0.2 17.2 0.04 30.1
HQLPP 33.9 0.01 66.5
@@
PGBRV0.2 14.1 0.01 27.7
HQLPP 19.0 0.01 35.8
BCC
KS GBRV 0.00 0.04 0.00
TF0.2vW GBRV 42.8 0.05 45.7
PGBRV0.2 20.7 0.04 25.4
HQLPP 34.7 0.00 49.3
@@
FCC
KS GBRV 0.00 -5490.58 0.00
TF0.2vW GBRV 81.0 -4613.32 63.5
PGBRV0.2 11.5 -6947.80 56.1
HQLPP 78.5 -5599.68 73.5
revHC GBRV 115.0 -4625.80 66.7
PGBRV0.2 27.2 6979.97 75.7
@@
PGBRV0.2 8.3 -0.038 56.7
HQLPP 84.4 -5483.58 52.8
BCC
KS GBRV 0.00 0.05 0.00
TF0.2vW GBRV 81.3 0.02 64.5
PGBRV0.2 11.1 0.01 55.9
HQLPP 116.3 0.00 98.4
revHC GBRV 144.0 0.07 86.6
PGBRV0.2 24.8 0.00 73.7
@@
FCC
KS GBRV 0.00 -2704.79 0.00
TF0.2vW GBRV 68.8 -2336.16 69.7
PGBRV0.2 14.5 -2929.42 61.8
HQLPP 55.2 -3544.12 76.3
revHC GBRV 51.6 -2343.41 51.3
PGBRV0.2 25.2 -2927.97 7.9
HQLPP 31.5 -3488.77 71.5
HCP
KS GBRV 0.00 0.03 0.00
TF0.2vW GBRV 58.8 0.20 84.6
PGBRV0.2 11.0 0.00 59.2
HQLPP 43.6 .01 68.9
revHC GBRV 45.2 0.061 66.3
PGBRV0.2 26.4 -0.006 13.9
@@
BCC
KS GBRV 0.00 0.05 0.00
TF0.2vW GBRV 67.7 0.02 67.3
PGBRV0.2 14.1 0.02 60.5
HQLPP 52.4 0.00 73.2
revHC GBRV 47.8 -0.08 43.6
PGBRV0.2 26.6 0.00 14.1
HQLPP 22.0 0.00 60.9
@@
BCC
KS GBRV 0.00 -1897.97 0.00
TF0.2vW GBRV 22.3 -2140.32 72.4
PGBRV0.2 38.1 -1990.97 74.8
HQLPP 40.9 -1906.13 77.3
revHC GBRV 3.2 -2145.08 47.9
PGBRV0.2 19.7 -1994.48 62.2
HQLPP 11.3 -1863.92 9.8
FCC
KS GBRV 0.00 0.43 0.00
TF0.2vW GBRV 20.4 -0.02 69.6
PGBRV0.2 34.7 0.04 74.5
HQLPP 38.4 0.00 74.1
revHC GBRV 2.3 0.11 45.6
PGBRV0.2 19.9 0.00 59.3
HQLPP 11.8 0.01 5.7
HCP
KS GBRV 0.00 0.44 0.00
TF0.2vW GBRV 17.2 0.01 69.6
PGBRV0.2 30.0 0.04 73.7
HQLPP 36.9 0.00 78.2
revHC GBRV 1.5 0.02 56.4
PGBRV0.2 18.5 0.0 65.7
HQLPP 12.1 0.0 23.4

### v3

@@ I Introduction and Background
In OF-DFT, the noninteracting kinetic energy, $T_{s}$, is evaluated approximately by a pure density functional, which we call henceforth the kinetic energy density functional or KEDF. Using KEDFs allows the energy minimization of the functional in Eq. (1) to be based solely on the electron density, completely bypassing the KS orbitals, and generally leading to algorithms of linear scaling complexity [mi2023orbital]. The low computational scaling resulting from the employment of approximate KEDFs comes at an accuracy cost. In fact OF-DFT development mostly focuses on development of KEDFs which are typically either semilocal [laricchia2011generalized, constantin2011semiclassical, luo2018simple] or nonlocal [huang2010nonlocal, shao2021revised, mi2019orbital]. In this work, we will not focus on KEDF development but we recognize that the accuracy of any result in this work is heavily biased by the quality of the KEDF employed.
@@
The guiding principles to build LPPs have been to employ them in a KS-DFT calculation (i.e., with exact $T_{s}$) imposing it to reproduce the electron density, lattice constant and even the band structure of solids [mi2016first, huang2008transferable, zhou2004transferable]. For those elements for which this is possible, it is said LPPs can be employed in OF-DFT calculations. It is certainly surprising that band structures can be reproduced in pseudopotential calculations having only a local part. However, in many cases it was shown to be possible [starkloff1977local]. It turns out that LPPs can be successfully constructed for the main-group elements mentioned above (provided that the semicore states are excluded from the pseudization [chi2024high]) and also for a large array of transition metals (where instead semicore electrons are included). The most accurate LPPs available to date for transition metal elements are the recently developed highly accurate LPPs by the Huang group [chi2024high]. The valence density of transition metals consists of contributions from both localized $d$ states and delocalized sspp states. Because existing KEDFs are based on the response of the homogeneous or nearly homogeneous electron gas, they struggle to provide accurate kinetic energies for transition metals [huang2012toward]. In addition, accurate pseudizations of transition metals should account for semicore states that lie close in energy to the valence $d$ bands and influence bonding and screening. Since OF-DFT represents $T_{s}[n]$ as a pure functional of the electron density, it lacks explicit angular-momentum dependence [ke2013angular] and therefore cannot distinguish $d$-state character from $s$ or $p$ contributions. These combined factors explain why transition metals are a particularly demanding class of atoms for OF-DFT and motivate the construction of pseudopotentials that mitigate these limitations.
@@ II.1 Nonlocal Pseudopotential Representations
Hohenberg and Kohn theorems for nonlocal external potentials due to Gilbert [gilbert1975hohenberg] apply and guarantee the existence of the above map, even though it is generally not bijective. The non-bijectivity is due to the fact that for a given choice of $\hat{v}_{nl}$ one can always find an associated $v_{loc}({\mathbf{r}})$ such that the KS equations yield the target density $n({\mathbf{r}})$. Therefore there are multiple (potentially an infinite number) of $\hat{v}_{NLPP}$ that have nonzero nonlocal part and that yield the same electron density, $n({\mathbf{r}})$. We call the set of such NLPPs a NLPP representation. Typically, there is one NLPP having the property that beyond $R_{cut}^{nl}$ not only the pseudodensity matches the all-electron density, but also the pseudo KS wavefunctions match the all-electron KS wavefunction for the isolated atom [kleinman1982efficacious].
@@ III Simultaneously approximating $\Delta v_{LPP}({\mathbf{r}})$ and the OF-DFT OEP problem
This definition clearly shows two goals for $\Delta v_{LPP}({\mathbf{r}})$: (1) it includes the effect of the nonlocal part of the KS-DFT PP; (2) it corrects for the fact that KEDFs are approximate, making sure that the density of the OF-DFT system matches the one of the KS-DFT system. Thus, the result of approximating $\Delta v_{LPP}({\mathbf{r}})$ is to simultaneously achieve the two goals.
@@
Important aspects to note are that $\Delta v_{LPP}({\mathbf{r}})$ is: (1) short ranged around each ion, as it is only nonzero for $0\leq|{\mathbf{r}}-\mathbf{R}_{I}|\leq r_{cut}$; (2) of spherical symmetry around each ion. Thus, the resulting LPP, given by $v_{loc}^{\hat{v}_{nl}}(r)+\Delta v_{LPP}(r)$ maintains the correct Coulomb decay in agreement with the target NLPP. Choosing an appropriate valence configuration for a pseudoatom is a complex task and has been addressed extensively in the pseudopotential literature. In this work, we simply adopt the valence configuration and the number of valence electrons from the corresponding target KS-DFT pseudopotential. As shown in Table S1, for some elements the number of valence pseudo-electrons (zval) differs between the two types of pseudopotentials considered. This difference arises from variations in zval among the corresponding KS-DFT pseudopotentials. We also note that there is freedom in choosing the cutoff radii ($r_{cut}$). We use the value of the largest of the $R_{cut}^{nl}$ of the target NLPP, see supplementary Table S1. We will see in the results section that this choice might not be the most optimal given the restrictive, polynomial ansatz for $\Delta v_{LPP}(r)$. The choice of cutoff radius will deserve further consideration in future works.
Regarding the choice of polynomial ansatz for $\Delta v_{LPP}(r)$, after testing, we found that a 5th order provides a good balance between accuracy of the LPPs and computational cost of the LPP optimizer. Reducing the order of the polynomial would result in significantly lower quality LPPs, and increasing the order would result is significant computational cost and convergence complexity for the optimizer.
Finally, we note that the proposed pseudization results in a overall smooth PP. The short-ranged polynomial is necessarily smooth because the high-order terms (beyond 5th) are not included. The Coulomb tail is recovered from the LPP of the target NLPP. Thus, the new LPPs are designed to be numerically well-behaved.
@@ IV Computational Details
All calculations employed the Local Density Approximation (LDA) xc functional proposed by Perdew and Zunger [perdew1981self]. Therefore, target results in this work are given by pseudopotential KS-DFT simulations employing the LDA xc functional. We also tested our pseudization procedure using PBE [perdew1996generalized] for Ag recovering a similarly accurate pseudization as LDA (see Results section). To test the effects that different xc functionals may have on the pseudization procedure, we included Tables S10-S12 collecting EOS parameters for Ag using PBE. Testing other xc functionals will be matter of future work.
@@ V.1 Generation of the new LPPs
where $N_{v}$ is the number of valence electrons. We note that $\Delta n\leq N_{v}$, thus $\Delta_{v}\leq 100$. In Supplementary Information Figure S1 we show the spatial distribution of the electron density difference between the OF and KS, using a representative cut plane and line cut for the most stable phases for Ag, Cu, and Pd. The figure validates the convergence criteria in Eq. (20).
@@ V.2 Equation of State
From the previous analysis, we have chosen to further investigate PGBRV0.2 for a selected set of atoms such as Cu, Ag, and Pd for which PGBRV0.2 is expected to improve upon the local part of the GBRV PP (i.e., the $v_{loc}^{\hat{v}_{nl}}$ mentioned before). We also present Mo for which we predict the PGBRV0.2 will not provide improved electronic structure compared to the local part of the GBRV NLPP. The analysis involves computation of EOS parameters (equilibrium volume, $V_{0}$; equilibrium energy, $E_{0}$; and bulk modulus, $B_{0}$) for FCC, BCC, and HCP phases. We compare the PGBRV0.2 results with the local part of the GBRV NLPP and HQLPP. All calculations are carried out with TF0.2vW and the nonlocal revHC KEDFs. The results are collected in Tables 1 for Ag, 2 for Cu, 3 for Pd, and 4 for Mo, where we show the relative error of $V_{0}$ and $B_{0}$ and the energy difference compared to the most stable phase of the presented elements. The EOS parameters from KS (LDA) are compared to experimental values in Table S9 showing that the KS-DFT benchmark results are close to experimental ground truth for lattice constants and within 30% for bulk moduli.
@@
Even though the results presented in Tables 14, as well as in supplementary Tables S2-S5, are not quantitatively accurate, they are nonetheless reasonable. In Table S6 we show a statistical summary of Tables S2-S5, where the Mean Absolute Relative Error (MARE) values obtained for $V_{0}$ and $B_{0}$ using the original GBRV and PSL pseudopotentials indicate high deviations between the OF-DFT and KS-DFT results. In contrast, the new local pseudopotentials (PGBRV0.2 and PPSL0.2) significantly reduce the MARE for $V_{0}$ to 28–37%, corresponding to a significant improvement in accuracy relative to employing the local part of the KS-DFT pseudopotentials. Although the $B_{0}$ values remain less accurate (MARE $\approx$55–60%), all materials are predicted to be bound. This results are nontrivial for OF-DFT calculations[rios2024effective, witt2021random], and as previously noted, the OF-DFT calculations employing the newly developed LPPs successfully capture several trends observed in the KS-DFT EOS parameters. To rationalize this outcome, we emphasize that the PGBRV0.2 LPP is designed not only to partially recover effects associated with the nonlocal component of the NLPP, but also, crucially, to locally invert the KS density in regions near the ionic cores. Consequently, using the PGBRV0.2 pseudopotentials yields electron densities of substantially higher quality than would be obtained using a generic, KEDF-agnostic LPP. A clear demonstration of this point is provided by the performance comparison with HQLPP pseudopotentials, which yield lower accuracy than PGBRV0.2. A more detailed quantitative comparison between the performance of PGBRV0.2 and HQLPP pseudopotentials is presented in Table S7, where statistical metrics are reported for the FCC, BCC, and HCP phases of Ag, Cu, Pd, and Mo using the TF0.2vW functional. When switching to the revHC functional (Table S8), the PGBRV0.2 pseudopotentials retain a similar level of accuracy (MARE $\approx$ 20–40 %), whereas the HQLPPs show noticeably larger deviations. These results confirm that the new pseudopotentials achieve improved consistency with respect to KS-DFT and additionally that their improved performance over using the local part of the KS-DFT PPs is retained across various KEDFs. This discrepancy does not reflect a flaw in the HQLPP pseudopotentials, as they were explicitly constructed to best represent electron densities calculated with the exact KEDF.
@@
FCC KS GBRV 0.00 -4023.32 0.00
TF0.2vW GBRV +43.0 -3719.80 -44.6
PGBRV0.2 -20.1 -4546.42 -22.3
HQLPP +34.7 -4054.91 -50.4
revHC GBRV +58.1 -3731.58 -56.8
PGBRV0.2 -13.3 -4380.85 -12.2
HQLPP +22.6 -3998.92 -50.4
HCP KS GBRV 0.00 0.01 0.00
TF0.2vW GBRV +35.8 0.14 -59.5
PGBRV0.2 -17.2 0.04 -30.1
HQLPP +33.9 0.01 -66.5
@@
PGBRV0.2 -14.1 0.01 -27.7
HQLPP +19.0 0.01 -35.8
BCC KS GBRV 0.00 0.04 0.00
TF0.2vW GBRV +42.8 0.05 -45.7
PGBRV0.2 -20.7 0.04 -25.4
HQLPP +34.7 0.00 -49.3
@@
FCC KS GBRV 0.00 -5490.58 0.00
TF0.2vW GBRV +81.0 -4613.32 -63.5
PGBRV0.2 +11.5 -6947.80 -56.1
HQLPP +78.5 -5599.68 -73.5
revHC GBRV +115.0 -4625.80 -66.7
PGBRV0.2 +27.2 6979.97 -75.7
@@
PGBRV0.2 -8.3 -0.038 -56.7
HQLPP +84.4 -5483.58 -52.8
BCC KS GBRV 0.00 0.05 0.00
TF0.2vW GBRV +81.3 0.02 -64.5
PGBRV0.2 +11.1 0.01 -55.9
HQLPP +116.3 0.00 -98.4
revHC GBRV +144.0 0.07 -86.6
PGBRV0.2 +24.8 0.00 -73.7
@@
FCC KS GBRV 0.00 -2704.79 0.00
TF0.2vW GBRV +68.8 -2336.16 -69.7
PGBRV0.2 +14.5 -2929.42 -61.8
HQLPP +55.2 -3544.12 -76.3
revHC GBRV +51.6 -2343.41 -51.3
PGBRV0.2 -25.2 -2927.97 +7.9
HQLPP +31.5 -3488.77 -71.5
HCP KS GBRV 0.00 0.03 0.00
TF0.2vW GBRV +58.8 0.20 -84.6
PGBRV0.2 +11.0 0.00 -59.2
HQLPP +43.6 .01 -68.9
revHC GBRV +45.2 0.061 -66.3
PGBRV0.2 -26.4 -0.006 -13.9
@@
BCC KS GBRV 0.00 0.05 0.00
TF0.2vW GBRV +67.7 0.02 -67.3
PGBRV0.2 +14.1 0.02 -60.5
HQLPP +52.4 0.00 -73.2
revHC GBRV +47.8 -0.08 -43.6
PGBRV0.2 -26.6 0.00 +14.1
HQLPP +22.0 0.00 -60.9
@@
BCC KS GBRV 0.00 -1897.97 0.00
TF0.2vW GBRV +22.3 -2140.32 -72.4
PGBRV0.2 +38.1 -1990.97 -74.8
HQLPP +40.9 -1906.13 -77.3
revHC GBRV -3.2 -2145.08 -47.9
PGBRV0.2 +19.7 -1994.48 -62.2
HQLPP -11.3 -1863.92 -9.8
FCC KS GBRV 0.00 0.43 0.00
TF0.2vW GBRV +20.4 -0.02 -69.6
PGBRV0.2 +34.7 0.04 -74.5
HQLPP +38.4 0.00 -74.1
revHC GBRV -2.3 0.11 -45.6
PGBRV0.2 +19.9 0.00 -59.3
HQLPP -11.8 0.01 -5.7
HCP KS GBRV 0.00 0.44 0.00
TF0.2vW GBRV +17.2 0.01 -69.6
PGBRV0.2 +30.0 0.04 -73.7
HQLPP +36.9 0.00 -78.2
revHC GBRV -1.5 0.02 -56.4
PGBRV0.2 +18.5 0.0 -65.7
HQLPP -12.1 0.0 -23.4